OpenStudy (anonymous):

What is the rate of change and initial value for the linear relation that includes the points shown in the table? x y 2 3 4 5 6 7 8 9 Initial value: 1, rate of change: 2 Initial value: 1, rate of change: 1 Initial value: 2, rate of change: 2 Initial value: 3, rate of change: 2

OpenStudy (anonymous):

@Michele_Laino caan you help? I have no Idea

OpenStudy (michele_laino):

initial value is the value of \(y\) at \(x=2\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its +1?

OpenStudy (anonymous):

So how would you explain the answer choices lol

OpenStudy (michele_laino):

please wait a moment, I'm thinking...

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

i think d

OpenStudy (michele_laino):

the data you provided can be modeled by this linear function: \(y=x+1\)

OpenStudy (anonymous):

mm hmm

OpenStudy (anonymous):

so d?

OpenStudy (michele_laino):

the rate of change \(r\), is: \[r = \frac{{7 - 5}}{{6 - 4}} = ...?\] or it is the slope of this equation: \(y=x+1\)

OpenStudy (anonymous):

oh. so 2, 2 right so like c

OpenStudy (michele_laino):

I have chosen the the second and third ordered pairs

OpenStudy (michele_laino):

I got this: \[r = \frac{{7 - 5}}{{6 - 4}} = 1\]

OpenStudy (anonymous):

lol ok so b

OpenStudy (michele_laino):

yes! I think it is b), since if we replace \(x=0\), we get: \(y=x+1=0+1=1\)

OpenStudy (anonymous):

lol thanks i have a few more could you help? ill repost

OpenStudy (michele_laino):

namely the initial value is the value of \(y\) at \(x=0\)

OpenStudy (anonymous):

:)

OpenStudy (michele_laino):

I can help! :)